Donald,Kate and maggie win £2400 between them on the national lottery.Donald gets a share of £x,whilst kate gets twice as much as donald.maggies share is £232 less than donalds amount.
write down the amount maggie and kate receive.

Answer by gonzo(654)

total amount = 2400 pounds (i think).
D gets x*2400
K gets 2*x*2400
M gets D - 232 = x*2400 - 232.
-----
x*2400 + 2*x*2400 + x*2400 - 232 = 2400 (equation 1)
-----
2400*x + 4800*x + 2400*x - 232 = 2400
9600*x - 232 = 2400
9600*x = 2400 + 232 = 2632
x = 2632 / 9600 = .274166667
2*x = .548333333
-----
substituting for x in equation 1 gets:
.274166667*2400 + 2*.274166667*2400 + .274166667*2400 - 232 = 2400
658 + 1316 + 658 - 232 = 2400
2400 = 2400.
answer is:
x = .274166667
donald gets 658
kate gets 1316
maggie gets 426